OpenStudy (anonymous):

The formation of disulfur is made from S8 and chlorine gas. the mixture of reactants are 27.5g of the sulfure and 35.2g of chlorine gas. what is the limiting reactant? Please show work so i can learn =)

OpenStudy (anonymous):

If we assume that the reaction is S8 + Cl2--> 4S2 + Cl2 then you need the molar ratio between S8 and Cl2 to be 1:1, meaning M(S8)=8x32=256 g/mol, and M(Cl2)=2x35.5=71 g/mol n=m/M-->n(S8)=27.5/256=0.107 and n(Cl)=35.2/71=0.496, and the limiting reactant is the one that is less i. e. S8 Note that I don't have a clue if this is the reaction, but this is the way to solve that kind of problems